OpenStudy (anonymous):

Find the radius of a sphere with a surface area of 1,236. (use 3.14 for pi) Answer 98.41 9.92 18.08 309

OpenStudy (anonymous):

A = 4*pi*r^2 1236 = 4*3.14*r^2 1236/(4*3.14) = r^2 98.41 = r^2 r = root98.41 r = 9.92 (approx)